A General Power bond carries a coupon rate of 8.3%, has 9 years until maturity, and sells at a yield to maturity of 7.3%. (Assume annual interest payments.)

a. What interest payments do bondholders receive each year?

Interest Payments:

b. At what price does the bond sell? (Do not round intermediate calculations. Round your answer to 2 decimal places.)

Price:

c. What will happen to the bond price if the yield to maturity falls to 6.3%? (Do not round intermediate calculations. Round your answer to 2 decimal places.)

Price will __ by __

d. If the yield to maturity falls to 6.3%, will the current yield be less, or more, than the yield to maturity?

More or Less?
Answer #1

a) assumes a face value of $1,000,interest will be 1000*8.3%=$83/year

b) Calculation of Bond price

Annual Interest 83

YTM-7.3%

Par Value & Redemption Value-$1,000

Present value of 9 year interest @7.3% rate=6.43*83=533.9(534)

Redemtion price(assumes)-$1,000*.530=530

so Bond Value=$1064

c) When YTM is 6.3%,

Price would be 6.71*83 +530=1087

So Price will be increased by $23

d) Current yield=current interest/current price

=83/1000=8.3%

YTM=6.3%

So Current yield is more than YTM